Finding the area of a rectangle when the length is v+3 and the width is 5, the answer is 5v + 15. I understand how I arrived at 15 but I don't understand the 5v part.
Thank you

Answer by mananth(16946)

length = v+3
width =5
Area A = L*W
A= 5*(v+3)
A= 5*v+5*3
A=5v+15